the length of a rectangular garden is 7 feet more than three times the width. the perimeter of the garden is 86 feet. find the area of the garden.
Answer #1

D then acco C b B A ? let A B C D be the Rectangular gardent and its coidth is b according to question, the length of gardenDate Page b = 3b +7 (by (i)) .. l = 3x9 +7 27+7 = 34 ft (6=9, by vo Now, we know that the formula for calculating area of Rec
Answer #2

SOLUTION :


Let length of garden be L and width be W .

So, L = 3W + 7 

Perimeter, P = 2L + 2W = 2(3W + 7) + 2W = 8W + 14 

=> 86 = 8W + 14

=> 8W = 86 - 14 = 72

=> W = 72/8 = 9 ft

=> L = 3 * 9 + 7 = 34 ft 


So, Area of the garden 

= 34 * 9

= 306 sq. ft. (ANSWER).

answered by: Tulsiram Garg